OpenStudy (anonymous):

factor completely 5x^3+20x^2+3x+12

hero (hero):

5x^2(x+4)+3(x+4) (x+4)(5x^2+3)

myininaya (myininaya):

gcf(5x^3,20x^2)=5x^2 gcf(3x,12)=3 so we have 5x^2(x+4)+3(x+4) gcd(5x^2(x+4),3(x+4))=x+4 (x+4)(5x^2+3)